# Bounds on detection of Bell correlations with entangled ultra-cold atoms in optical lattices under occupation defects

### v1

@@ VII Summary and conclusions
Our proposed toy model accounts for imperfections that provide non-unit filling per site in the preparation and measurement stages through the probability $p$ of a site being single-occupied. The $M$-site Bell correlator shows a violation for this model in the limit of a large system, when $p>p_{c}\approx 1/\sqrt{2}$. On the other hand, the two-site Bell correlator relying on collective spin measurements shows the bounds $p_{c}=4/5$ and $p_{c}=\sqrt{3}/2$ when $N=pM<M$ and $N=M$, respectively. We illustrate these general results under two practical methods of entanglement generation varied by the source of imperfections: vacant sites due to initial non unit filling, and vacant and multiply occupied sites due to non-zero temperature of the initial state. In the presence of holes, we study the generation of entanglement in the Mott insulating via interaction anisotropy and inhomogeneous field demonstrating the validity of the toy model predictions for the critical value of the filling factor allowing generation of Bell correlations. For non-zero temperatures, we study the generation of entanglement in the superfluid regime produced via atom-atom interactions transferred to the Mott regime via adiabatic rising of the lattice height. We find a connection between the probability $p$ of the toy model and the effective temperature $T$ of the initial state that drives the system from the unit filling regime under study. We identified the critical value of the initial temperature allowing violation of Bell inequalities.
Our results reveal the fundamental limits on detecting Bell correlations in the lattice system due to occupation defects. The bounds identified in specific protocols fall within the range of typical experimental realizations, suggesting the feasibility of detecting Bell correlations.

### v2

@@ VII Summary and conclusions
Our proposed toy model accounts for imperfections that provide non-unit filling per site in the preparation and measurement stages through the probability $p$ of a site being single-occupied. The $M$-site correlator shows a violation of the Bell inequality for GHZ states when $p>p_{c}=2^{1/M}/\sqrt{2}$. On the other hand, the two-site Bell correlator relying on collective spin measurements for spin-squeezed states shows the bounds $p_{c}=4/5$ when $N=pM<M$, and $p_{c}=\sqrt{3}/2$ when $N=M$. We illustrate these general results using two different methods of entanglement generation implying different sources of imperfections: vacant sites due to initial non-unit filling, and vacant and multiply occupied sites due to non-zero temperature of the initial state. In the presence of holes, we study the generation of entanglement in the Mott insulating via interaction anisotropy and inhomogeneous field demonstrating the validity of the toy model predictions for the critical value of the filling factor allowing the detection of violation of Bell inequalities. For non-zero temperatures, we study the entanglement generation in the superfluid regime produced via atom-atom interactions transferred to the Mott regime via adiabatic rising of the lattice height. We find a connection between the probability $p$, of a single occupation of a given site, of the toy model and the effective temperature $T$ of the initial state. We then identified the critical value of the initial temperature allowing violation of Bell inequalities.
Our results reveal the fundamental limits on detecting Bell correlations in the lattice system due to occupation defects. Analytical predictions of the toy model are general and can be relevant for any platform where addressing individual spins is possible. The bounds identified in specific protocols fall within the range of typical experimental realizations, suggesting the feasibility of detecting Bell correlations.